CAT Chat Minutes

June 28, 2004

 

Information and Follow-up
As a viewgraph of the current run schedule was projected, user noted that there appears to be no regular pattern to the scheduling of machine studies between 48 hr. intervention periods.   In preparing schedules for the future, can study periods be more regular?

Roger explained that the schedule is shifted around to accommodate the holidays, and is subject to the 5000 operations hour requirement for the fiscal year.   We'll look into the possibility of having the short interventions occurring on a more regular pattern.

General Information
Machine Stats
 
Delivered beam:   673.64
   Percentage of scheduled time:   97.90%
   Downtime during period:   14.45 hrs.
   Mean time between faults:   44.91 hrs.

Faults:

# 22   06/23 08:00   To   06/23 13:19     5.33  

Power bump  [ComEd ]

  0.93 CA problem, bunch purity tuning, refill

# 23   06/23 14:15   To   06/26 05:16    63.02  

Valve ctl failure   [CTL]

  3.60 Valve ctl PS replaced 2.2, S2 absorber flow 1.4 ME

# 25   06/26 08:52   To   06/28 08:04    47.21  

Fill Ongoing

  0.00  

Beamline downtime summary by APS groups, 06/21/2004 to 06/28/2004

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

410

06/10/04

18-ID

Open

28 min

Mono shutter located in station C faulted shortly after opening; PSS fault code #211 P8DMS2_NOSWITCH.

Problem continues. Resolution to be determined.

410

06/10/04

18-ID

Open

5 min

ASD-VAC

Mono shutter located in station C faulted shortly after opening; PSS fault code #211 P8DMS2_NOSWITCH.

Problem continues. Resolution to be determined.

419

06/23/04

33-BM

Closed

54 min

ASD-SI

FE-EPS trip reported by Beamline Personnel. EPICS Vacuum screen (S_BMvac.adl) indicated that IP8 caused the trip.

ASD-SI modified FE-EPS to require redundant confirmation of "bad" vacuum signals from IP7 and IP8 before issuing close command to PS2. Modification to FE-EPS is temporary for the duration of run 2004-2. Problem to be mitigated during next shutdown.

Machine studies planned for 4:00 p.m. today.

APS Spectrum Picnic:   Because of the great numbers attending the barbeque lunch, picnic tables from LOMs are being located to the gallery patio for use on Wednesday, June 30 th.   All picnic tables will be returned following the lunch.